OpenStudy (anonymous):

In an experiment, 400. mL of 0.600 M HNO3(aq) is mixed with 400. mL of 0.300 M Ba(OH)2(aq) in a constant-pressure calorimeter having a heat capacity of 387 J/oC. The initial temperature of both solutions is the same at 18.88oC, and the final temperature of the mixed solution is 22.49oC. Calculate the heat of neutralization in kJ per mole of HNO3.

OpenStudy (aaronq):

Find the heat produced, q, with \(q=m*C*\Delta T\); where: m=mass (in this instance, of the solution, assume the density of water d=1 g/mL), C=specific heat capacity, \(\Delta T=T_f-T_i\) is the change in temperature. Use: \(Molarity=\dfrac{n_{HNO_3}}{L_{solution}}\) to find moles of HNO3, the divide q by the moles. \(\dfrac{q_{produced}}{n_{HNO_3}}=heat ~per ~mole\)

OpenStudy (anonymous):

thanks i just wanted to make sure i was taking the right steps.